OpenStudy (anonymous):

If 1100 square centimeters of material is available to make a box with a square base and an open top, find the largest possible volume of the box.

OpenStudy (anonymous):

got a picture?

OpenStudy (anonymous):

if you call the lenght of the base \(x\) and the height \(h\) then the base has area \(x^2\) and the four sides have area \(xh\) making the total \[x^2+4xh\]

OpenStudy (anonymous):

since you know that total is \(1100\) you can set \[x^2+4xh=1100\] and solve for \(h\) in terms of \(x\)

OpenStudy (anonymous):

pretty much in your head i get \[h=\frac{1100-x^2}{x}\]

OpenStudy (anonymous):

then the volume is \[V=x^2h\] but now it is \[V(x)=x^2(\frac{1100-x^2}{x})\] maximize that one

OpenStudy (anonymous):

you good from there?